Jordan buys a bag of cookies that contains 4 chocolate chip cookies, 4 peanut butter cookies, 8 sugar cookies and 8 oatmeal cookies. What is the probability that Jordan randomly selects a chocolate chip cookie from the bag, eats it, then randomly selects a sugar cookie? Express your answer as a reduced fraction.

Found 2 solutions by greenestamps, JmeLs:
Answer by greenestamps(13200)



24 cookies initially, of which 4 are chocolate chip: P=4/24

Then 23 cookies remaining, of which 8 are sugar: P=8/23

P(chocolate chip then sugar) = (4/24)(8/23)

Perform the multiplication and simplify or convert to decimal as desired/required.


Answer by JmeLs(1)